# New simple and accurate approximations for quintessence

### v1

@@ 3.2 Tracking quintessence
The dynamical system (51) admits $\tilde{\phi}=0$ (and thereby $\phi=0$) as an invariant boundary subset. As shown in [31], tracking quintessence is associated with a tracker orbit that originates from a tracker fixed point, $\mathrm{T}$, on the $\tilde{\phi}=0$ boundary, which in the present variables is given by
@@ 4 Numerical comparisons
In this section we calculate numerical relative errors, i.e.,
@@
$\Omega_{\mathrm{DE}}$ $\approx\frac{\Omega_{\mathrm{DE},0}}{\Omega_{\mathrm{m}0}(1+z)^{-3w_{\infty}}\left(\frac{1+\beta T}{1+\beta T_{0}}\right)^{\frac{\gamma-1}{\beta}}+\Omega_{\mathrm{DE},0}},$ (59b)
$E_{\mathrm{DE}}^{2}$ $\approx\Omega_{\mathrm{m}0}(1+z)^{3}+\Omega_{\mathrm{DE}0}(1+z)^{3(1+w_{\infty})}\left(\frac{1+\beta T_{0}}{1+\beta T}\right)^{\frac{\gamma-1}{\beta}},$ (59c)
@@
For thawing quintessence we will consider a currently popular model, the quintessential $\alpha$-attractor inflation model introduced in [47] and further explored in [48], [49] and [38], referred to as Exp-model I in [48], [49] and as the EC potential in [38], given by
@@
To obtain an explicit model, we set the model parameters $\alpha$ and $\nu$ as follows: $(\alpha,\nu)=(\frac{7}{3},128)$. We considered these models in [38] in the context quintessential inflation, where inflationary considerations determined $V_{-}$, see [48] and [49]. The inclusion of radiation yielded an integral, which resulted in that only one of the thawing solutions in the 1-parameter set characterized by $\phi_{*}$ was compatible with the chosen values of $V_{-}$, $H_{0}$, $\Omega_{\gamma 0}$ and $\Omega_{\mathrm{m}0}$. This solution was found by varying the frozen scalar field value, $\phi_{*}$, at a line of fixed points from which the thawing solutions originated, which resulted in $\phi_{*}=9.594$.
In the present case where we neglect radiation the integral does not exist and then $H_{0}$ and $\Omega_{\phi 0}=0.68$ do not impose any restrictions on $\phi_{*}$ at the line of fixed points $(\phi,\Sigma_{-},\Omega_{\phi})=(\phi_{*},0,0)$ from which the thawing quintessence oribits originate. However, due to that radiation is negligible when quintessence affect the evolution, the present approximations are excellent also when radiation is included as in [38]. For this reason we choose $\phi_{*}=9.594$. Inserting this value into (61) enables the subsequent computation of $\gamma$ and $\beta$ by means of (45) and (46), which results in $(\gamma,\beta)=(1.095,0.789)$, thereby making the approximation explicit. To obtain the relative errors we numerically integrate the dynamical system (38) and the decoupled equation for $E_{\phi}$,
@@ 5 Discussion
  • Instead of using $w_{\mathrm{DE}}$ as the starting point, one can use other variables. For example, as pointed out in footnote 7, our series expansions also give rise to a series expansion in $T$ for $E_{\mathrm{DE}}^{2}$ and thereby also for $E_{\mathrm{DE}}$. This expansion can subsequently be used to obtain various Padé approximants, which can then serve as the starting point for calculating various other quantities. Or one can derive series expansions for one of the distance measures and use this for constructing Padé approximants, from which other variables and observables can be derived.
  • We have here neglected radiation, which, however, can be included in several different ways; one can e.g. use the fixed points on the radiation boundary instead of the present matter boundary when it comes to quintessence, as was done in [38]. Or one can simply add radiation, $\rho_{\gamma}=\rho_{\gamma 0}\exp(-4N)=3H_{0}^{2}\Omega_{\gamma 0}\exp(-4N)$, to $\rho$, i.e. $\rho=\rho_{\gamma}+\rho_{\mathrm{m}}+\rho_{\mathrm{DE}}$, and keep the present $w_{\mathrm{DE}}$ expressions, since radiation has a negligible influence on quintessence evolution, and, e.g., use that $\rho=3H^{2}$.
@@
  • One can use other dynamical systems variables than the present ones to obtain somewhat different results. For example, we could have used the variables in [31] to deal with tracking quintessence, which would have been useful for the hyperbolic potentials that in addition to the inverse power law potential was dealt with in that paper. [^17] This would had come with the price that $\Gamma$ would not have appeared explicitly in the results for $\gamma$ and $\beta$ due to the different asymptotic regularization of $\lambda$. However, to improve convergence one should then replace the scalar field variable $\bar{\varphi}$ in [31] with $\tilde{\varphi}=\bar{\varphi}^{2}$.
@@
  • We have here focussed on obtaining thawing and tracking quintessence approximations situated in a broader DE context for spatially homogeneous, isotropic and flat spacetimes. A natural next step is to use these approximations in order to interpret the rapidly growing set of increasingly precise observational data. In particular, the present results, or variations thereof, can serve as input in cosmological perturbation theory to bring the results closer to a wide set of observations.
  • We finally note that the present approach and variations thereof can easily be applied to other models and theories, especially for those admitting an Einstein frame, although this usually adds parameters thereby further diminishing predictive power.

### v2

@@ 3.2 Tracking quintessence
The dynamical system (52) admits $\tilde{\phi}=0$ (and thereby $\phi=0$) as an invariant boundary subset. As shown in [32], in regularizing variables tracking quintessence is associated with a tracker orbit that originates from a tracker fixed point, $\mathrm{T}$, on the $\tilde{\phi}=0$ boundary, which in the present variables is given by
@@ 4 Numerical comparisons
In this section we calculate the numerical relative errors of the approximations, i.e.,
@@
$\Omega_{\mathrm{DE}}$ $\approx\frac{\Omega_{\mathrm{DE}0}}{\Omega_{\mathrm{m}0}(1+z)^{-3w_{\infty}}\left(\frac{1+\beta T}{1+\beta T_{0}}\right)^{\frac{\gamma-1}{\beta}}+\Omega_{\mathrm{DE}0}},$ (60b)
$E_{\mathrm{DE}}^{2}$ $\approx(1+z)^{3}\left[\Omega_{\mathrm{m}0}+\Omega_{\mathrm{DE}0}(1+z)^{3w_{\infty}}\left(\frac{1+\beta T_{0}}{1+\beta T}\right)^{\frac{\gamma-1}{\beta}}\right],$ (60c)
@@
For thawing quintessence we will consider a currently popular model, the quintessential $\alpha$-attractor inflation model introduced in [49] and further explored in [50], [51] and [40], referred to as Exp-model I in [50], [51] and as the EC potential model in [40], given by
@@
To obtain an explicit model, we set the model parameters $\alpha$ and $\nu$ as follows: $(\alpha,\nu)=(\frac{7}{3},128)$. We considered these models in [40] in the context quintessential inflation, where inflationary considerations determined $V_{-}$, see [50] and [51]. The inclusion of radiation yielded an integral, which resulted in that only one of the thawing solutions in the 1-parameter set characterized by $\phi_{*}$ was compatible with the chosen values of $V_{-}$, $H_{0}$, $\Omega_{\gamma 0}$ and $\Omega_{\mathrm{m}0}$. This solution was found by varying the frozen scalar field value, $\phi_{*}$, at the line of fixed points from which the thawing solutions originated, which resulted in $\phi_{*}=9.594$.
In the present case where we neglect radiation the integral does not exist and then $H_{0}$ and $\Omega_{\phi 0}=0.68$ do not impose any restrictions on $\phi_{*}$ at the line of fixed points $(\phi,\Sigma_{-},\Omega_{\phi})=(\phi_{*},0,0)$ from which the thawing quintessence oribits originate. However, due to that radiation is negligible when quintessence affects the evolution, the present approximations are excellent also when radiation is included as in [40]. For this reason we choose $\phi_{*}=9.594$. Inserting this value into (62) enables the subsequent computation of $\gamma$ and $\beta$ by means of (46) and (47), which results in $(\gamma,\beta)=(1.095,0.789)$, thereby making the approximation numerically explicit. To obtain the relative errors we numerically integrate the dynamical system (39) and the decoupled equation for $E_{\phi}$,
@@ 5 Discussion
  • Instead of using $w_{\mathrm{DE}}$ as the starting point, one can use other variables. For example, as pointed out in footnote 7, our series expansions also give rise to a series expansion in $T$ for $E_{\mathrm{DE}}^{2}$ and thereby also for $E_{\mathrm{DE}}$. This expansion can subsequently be used to obtain various Padé approximants, which can then serve as the starting point for calculating other quantities. Or one can derive series expansions for one of the distance measures and use this for constructing Padé approximants, from which other variables and observables can be derived.
  • We have here neglected radiation, which, however, can be included in several different ways; one can e.g. use the fixed points on the radiation boundary as the starting point instead of the fixed points on the matter boundary when it comes to quintessence, as was done in [40]. Or one can simply add radiation, $\rho_{\gamma}=\rho_{\gamma 0}\exp(-4N)=3H_{0}^{2}\Omega_{\gamma 0}\exp(-4N)$, to $\rho$, i.e. $\rho=\rho_{\gamma}+\rho_{\mathrm{m}}+\rho_{\mathrm{DE}}$, and keep the present $w_{\mathrm{DE}}$ expressions, since radiation has a negligible influence on quintessence evolution, and, e.g., use that $3H^{2}=\rho$.
@@
  • One can use other dynamical systems variables than the present ones to obtain somewhat different results. For example, we could have used the variables in [32] to deal with tracking quintessence, which would have been useful for the hyperbolic potentials that in addition to the inverse power law potential was dealt with in that paper. [^18] This would had come with the price that $\Gamma$ would not have appeared explicitly in the results for $\gamma$ and $\beta$ due to the different regularization of $\lambda$. However, to improve convergence one should then replace the scalar field variable $\bar{\varphi}$ in [32] with $\tilde{\varphi}=\bar{\varphi}^{2}$.
@@
  • We have here focussed on obtaining thawing and tracking quintessence approximations situated in a broader DE context for spatially homogeneous, isotropic and flat spacetimes. A natural next step is to use these approximations in order to interpret and assess the rapidly growing set of increasingly precise observational data. In particular, the present results, or variations thereof, can serve as input in cosmological perturbation theory to bring the results closer to a wide set of observations.
  • We finally note that the present approach and variations thereof can easily be applied to other models and theories, especially for those admitting an Einstein frame, although this usually adds parameters that diminish predictive power.